# Estimates of Kähler metrics on noncompact finite volume hyperbolic Riemann surfaces, and their symmetric products

### v2

@@ 1.2. Statements of main results
Since the Bergman kernel $\operatorname{\mathcal{B}_{\Omega_{\it{\overline{X}}}}^{\it{k}}}$ vanishes at $i\infty$, the Bergman metric $\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}$ is not defined at $i\infty$. However, as shown in Proposition 2.8, the ratio
@@
where the implied constant depends only on $\operatorname{\Gamma}$.
@@
With notation as above, for a given $d\geq 1$, and $k\gg 1$, and $z\in\operatorname{\mathrm{Sym}^{{\it d}}({\it\overline{X}})}$, we have the following estimate
(3) $\sup_{z\in\operatorname{{\it{\overline{X}}}}}\bigg{|}\frac{\mu_{{\rm Sym}(\operatorname{{\it{\overline{X}}}}),\mathrm{vol}}^{\mathrm{FS},k}(z)}{\mu_{\operatorname{{\it{\overline{X}}}}^{d},\mathrm{vol}}(z)}\bigg{|}=O_{\operatorname{\Gamma}}(k^{3d}),$
where the implied constant depends only on $\operatorname{\Gamma}$.
@@
Furthermore, we apply the Fourier expansions of cusp forms in the neighborhood of $i\infty$, demonstrating the interplay between geometry and number theory. We refer the reader to Proposition 2.7 for an estimate on Fourier coefficients of cusp forms, which plays an important role in the proof of Main Theorem . Proposition 2.7 was proved by Kowalski in an unpublished manuscript [13]. We give an alternate proof using estimates of Bergman kernel.
@@ 2.1. Notation and background material
We now define the Fourier expansion of cusp forms. For $k\geq 1$, and $z\in\operatorname{\mathbb{H}}$, and $f\in\operatorname{\mathcal{S}^{2{\it{k}}}(\Gamma)}$ admits the following Fourier expansion, which is of the following form
[…]
The constants $\big{\{}a_{m}\big{\}}_{m\geq 1}$ are known as Fourier coefficients of $f$, and encode a lot of arithmetic and geometric information.
For a given $k\geq 1$, set
(14) $U_{\operatorname{\Gamma},k}:=\bigg{\{}z=x+iy\in{\mathcal{F}}_{\operatorname{\Gamma}}\bigg{|}\,y\geq\frac{k}{2\pi}\bigg{\}};\,\,\overline{U}_{\operatorname{\Gamma},k}:=U_{\operatorname{\Gamma},k}\cup\{i\infty\},\,\,\,{\mathcal{F}}_{\operatorname{\Gamma},k}:={\mathcal{F}}_{\operatorname{\Gamma}}\backslash U_{\operatorname{\Gamma},k}.$
For $f_{j}\in\operatorname{\mathcal{S}^{2{\it{k}}}(\Gamma)}$, let $\big{\{}a_{j,m}\big{\}}_{m\geq 1}$ denote the set of Fourier coefficients of the cusp form $f_{j}$, which are as described in equation (2.1).
@@ 2.2. Auxiliary estimates
From the proof of Theorem 6.1 in [10], we have
$\sup_{z\in\operatorname{{\it{\overline{X}}}}}\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=\sup_{z\in{\mathcal{F}}_{\operatorname{\Gamma},k}}\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}.$
Combining estimate (15) with the above equality completes the proof of the proposition. ∎
@@ 2.3. Estimates of the Bergman metric
So employing the same arguments as in the proof of Theorem 6.1 from [10], we can conclude that
@@
We now estimate the Bergman metric in the compact subset ${\mathcal{F}}_{\operatorname{\Gamma},k}$, which is as defined in equation (14).
Proposition 2.6.
@@
$\sup_{z\in{\mathcal{F}}_{\operatorname{\Gamma},k}}\bigg{|}\frac{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}(z)}{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{hyp}}}(z)}\bigg{|}=O_{\operatorname{\Gamma}}(k^{3}),$
@@
For $k\geq 3$, and $z\in{\mathcal{F}}_{\operatorname{\Gamma},k}$, from Lemma 2.3, we have
@@
We now estimate the second term on the right-hand side of the above inequality. For $z\in{\mathcal{F}}_{\operatorname{\Gamma},k}$, from estimate (22), we derive
@@
Furthermore, for $k\gg 1$, from the proof of Proposition 2.2, it is clear that
(25) $\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}\gg_{\operatorname{\Gamma}}k,$
@@
We now estimate the third term on the right-hand side of inequality (23). For $z\in{\mathcal{F}}_{\operatorname{\Gamma},k}$, combining estimates (21) and (25), we arrive at the following estimate, for the third term on the right-hand side of inequality (23)
@@
Combining equation (23) with estimates (26) and (27), completes the proof of the proposition. ∎
In order to estimate the Bergman metric in $\overline{U}_{\operatorname{\Gamma},k}$, the neighborhood of the puncture $i\infty$, we need to derive estimates of the Fourier coefficients of cusp forms, which we establish in the following proposition. The proposition is already proved by Kowalski in an unpublished manuscript [13]. However, for the benefit of the reader we give an alternate paper using the Bergman kernel.
Proposition 2.7.
With notation as above, for a fixed $k\gg 1$, let
${\mathcal{S}}:=\big{\{}f_{1},\ldots,f_{n_{k}}\big{\}}$
denote an orthonormal basis for $\operatorname{\mathcal{S}^{2{\it{k}}}(\Gamma)}$, with respect to the Petersson inner-product (which is as defined in equation (9)), and let $\{a_{j,n}\}_{n\geq 1}$ denote the set of Fourier coefficients of a given cusp form $f_{j}\in{\mathcal{S}}$. Then, for any $y\geq k/2\pi$, we have the following lower bound
(28) $\sum_{j=1}^{n_{k}}\sum_{n=1}^{\infty}\big{|}a_{j,n}\big{|}^{2}e^{-4\pi ny}\gg_{\operatorname{\Gamma}}\frac{k}{y^{2k}},$
Proof.
From the Fourier expansion of cusp forms (which is as described in equation (2.1)), we observe that
[…]
Using the series expansion of the Bergman kernel (11), and employing arguments from the proof of Proposition 2.2 we now compute
[…]
$\frac{2k-1}{4\pi}\bigg{|}\sum_{n\in{\mathbb{Z}}}\frac{(2i)^{2k}}{\big{(}2iy-n\big{)}^{2k}}\bigg{|}-\frac{2k-1}{4\pi y^{2k}}\bigg{(}\frac{16}{{\rm cosh}^{2k-4}(\operatorname{{\it{r}}_{\operatorname{\Gamma}}}/4)}+\frac{8}{{\rm cosh}^{2k-3}(\operatorname{{\it{r}}_{\operatorname{\Gamma}}}/2)}\bigg{)}-$
(30) $\frac{2k-1}{2\pi{\rm sinh}^{2}(\operatorname{{\it{r}}_{\operatorname{\Gamma}}}/4)y^{2k}}\bigg{(}\frac{1}{{\rm cosh}^{2k-3}(\operatorname{{\it{r}}_{\operatorname{\Gamma}}}/2)}+\frac{1}{{\rm cosh}^{2k-4}(\operatorname{{\it{r}}_{\operatorname{\Gamma}}}/2)}\bigg{)}.$
Realizing that
$\sum_{n\in{\mathbb{Z}}}\frac{1}{\big{(}2iy-n\big{)}^{2k}}=\zeta(2iy,k)+\zeta_{1}(-2iy,k),$
where $\zeta$ and $\zeta_{1}$ denote the Riemann-Hurwitz function and modified Riemann-Hurwitz function, respectively, and applying the well-known asymptotics of both the functions, for $y\geq k/2\pi$, and $k\gg 1$, we derive
(31) $\frac{2k-1}{4\pi}\bigg{|}\sum_{n\in{\mathbb{Z}}}\frac{(2i)^{2k}}{\big{(}2iy-n\big{)}^{2k}}\bigg{|}\gg\frac{k}{y^{2k}}+\frac{k}{(k-1)y^{2k-1}}\gg\frac{k}{y^{2k}}.$
Combining estimates (29)–(31), completes the proof of the lemma. ∎
Proposition 2.8.
With notation as above, for $k\gg 1$, we have the following estimate
$\sup_{z\in U_{\operatorname{\Gamma},k}}\bigg{|}\frac{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}(z)}{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{hyp}}}(z)}\bigg{|}=O_{\operatorname{\Gamma}}(k^{3}),$
where the the implied constant depends only on the Fuchsian subgroup $\operatorname{\Gamma}$.
Proof.
For $k\geq 2$, and $z\in U_{\operatorname{\Gamma},k}$, from Lemma 2.3, we have the following relation
(32) $\frac{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}(z)}{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{hyp}}}(z)}=\frac{k}{2\pi}+\frac{y^{2}}{\pi}\displaystyle\frac{\displaystyle\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z}\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial\overline{z}}}{\big{(}\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\big{)}^{2}}-\frac{y^{2}}{\pi}\displaystyle\frac{\displaystyle\frac{\partial^{2}\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z\partial\overline{z}}}{\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}.$
We now estimate the the second and third terms on the right-hand side of the above equation. For $k\gg 1$, and $z\in U_{\operatorname{\Gamma},k}$, using the Fourier expansion of cusp forms, which is as described in equation (2.1), and Proposition 2.7, we arrive at the following $q$-expansions
$\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)=\alpha(z)+\beta(z),$
$\mathrm{where}\,\,\alpha(z):=\sum_{j=1}^{n_{k}}\sum_{n=1}^{\infty}|a_{j,n}|^{2}\big{|}q(z)\big{|}^{2n}\gg_{\operatorname{\Gamma}}\frac{k}{y^{2k}},$
(33) $\mathrm{and}\,\,\beta(z):=\sum_{\begin{subarray}{c}m,n=1\\ m\not=n\end{subarray}}^{\infty}a_{j,n}\overline{a_{j,m}}q(nz)\overline{q(mz)}=O_{\operatorname{\Gamma}}\big{(}|q(z)|\big{)},$
and $\{a_{j,n}\}_{n\geq 1}$ denotes the set of Fourier coefficients of the cusp form $f_{j}\in\operatorname{\mathcal{S}^{2{\it{k}}}(\Gamma)}$, and $q(z)$ is as defined in equation (4), and $n_{k}$ is the dimension of $\operatorname{\mathcal{S}^{2{\it{k}}}(\Gamma)}$.
Combining equation (2.8) with estimates (28) and (22), we obtain the following estimate
$\frac{y^{2}}{\pi}\Bigg{|}\displaystyle\frac{\displaystyle\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z}\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial\overline{z}}}{\big{(}\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\big{)}^{2}}\Bigg{|}\leq\frac{y^{2}\bigg{|}\displaystyle\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z}\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial\overline{z}}\bigg{|}}{\pi|\alpha(z)|^{2}\cdot\bigg{|}1+\displaystyle\frac{\beta(z)}{\alpha(z)}\bigg{|}^{2}}\ll_{\operatorname{\Gamma}}$
(34) $\frac{y^{4k+2}\bigg{|}\displaystyle\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z}\frac{\partial\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial\overline{z}}\bigg{|}}{k^{2}\bigg{|}1+\displaystyle\frac{\beta(z)}{\alpha(z)}\bigg{|}^{2}}=O_{\operatorname{\Gamma}}(k^{3}).$
Similarly combining equation (2.8) awith estimates (28) and (21), we arrive at the following estimate
(35) $\frac{y^{2}}{\pi}\Bigg{|}\frac{\displaystyle\frac{\partial^{2}\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z\partial\overline{z}}}{\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}\Bigg{|}\leq\frac{y^{2}\bigg{|}\displaystyle\frac{\partial^{2}\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z\partial\overline{z}}\bigg{|}}{\pi|\alpha(z)|\cdot\bigg{|}1+\displaystyle\frac{\beta(z)}{\alpha(z)}\bigg{|}}\ll_{\operatorname{\Gamma}}\frac{y^{2k+2}\bigg{|}\displaystyle\frac{\partial^{2}\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)}{\partial z\partial\overline{z}}\bigg{|}}{k\bigg{|}1+\displaystyle\frac{\beta(z)}{\alpha(z)}\bigg{|}}=O_{\operatorname{\Gamma}}(k^{5/2}).$
Combining equation (32) with estimates (2.8)–(35), completes the proof of the proposition. ∎
Remark.
For a fixed $k\geq 3$, from Proposition 2.8, it is clear that the Bergman metric $\operatorname{\mu_{{\it X}}^{\mathrm{Ber},{\it{k}}}}(z)$ is not defined at the point $i\infty$. However, the ratio of the Bergman metric and hyperbolic metric remains bounded, as $z\in\operatorname{{\it{\overline{X}}}}\backslash\{i\infty\}$, approaches the puncture $i\infty$.
Combining Lemma 2.3 with Propositions 2.6, and 2.8 we now estimate the Bergman metric on $\operatorname{{\it{\overline{X}}}}$ in the following theorem, and hence, complete the proof of Main Theorem 1.
Theorem 2.9.
(36) $\bigg{|}\frac{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}(z)}{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{hyp}}}(z)}\bigg{|}=O_{\operatorname{\Gamma}}(k^{3}),$
Proof.
The proof of the theorem follows from combining Propositions 2.6 and 2.8. ∎

### v3

@@ 1.2. Statements of main results
Since the Bergman kernel $\operatorname{\mathcal{B}_{\Omega_{\it{\overline{X}}}}^{\it{k}}}$ vanishes at $i\infty$, the Bergman metric $\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}$ is not defined at $i\infty$. However, as shown in Theorem 2.7, the ratio
@@
With notation as above, for a given $d\geq 1$, and $k\gg 1$, we have the following estimate
(3) $\sup_{z\in\operatorname{\mathrm{Sym}^{{\it d}}({\it\overline{X}})}}\bigg{|}\frac{\mu_{{\rm Sym}(\operatorname{{\it{\overline{X}}}}),\mathrm{vol}}^{\mathrm{FS},k}(z)}{\mu_{\operatorname{{\it{\overline{X}}}}^{d},\mathrm{vol}}(z)}\bigg{|}=O_{\operatorname{\Gamma}}(k^{3d}),$
@@ 2.2. Auxiliary estimates
In the following proposition, combining arguments from Main Theorem in [5], and Theorem 6.1 in [10], we derive an estimate for the Bergman kernel. For the benefit of the reader, we reproduce the arguments here.
@@
We now apply the arguments from the proof of Theorem 6.1 in [10]. For a given $\varepsilon>0$, set
$U_{\operatorname{\Gamma},\varepsilon}:=\big{\{}z=x+iy\in{\mathcal{F}}\big{|}\,y\geq 1/\varepsilon\big{\}}\,\,\mathrm{and}\,\,{\mathcal{F}}_{\operatorname{\Gamma},\varepsilon}:={\mathcal{F}}\backslash U_{\operatorname{\Gamma},\varepsilon}.$
From estimate (13), it is clear that
(17) $\sup_{z\in{\mathcal{F}}_{\operatorname{\Gamma},\varepsilon}}\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=O_{\operatorname{\Gamma}}\big{(}k/\sqrt{\varepsilon}\big{)}.$
Writing
$\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=g(z)h(z),$
(18) $\mathrm{where}\,\,g(z):=y^{2k}|q(z)|^{2},\,\,\mathrm{and}\,\,h(z)=y^{2k}\sum_{j=1}^{n_{k}}\frac{|f_{j}(z)|^{2}}{|q(z)|^{2}},$
and observing that the function $h(z)$ is a subharmonic function, we deduce that
(19) $\sup_{z\in U_{\operatorname{\Gamma},\varepsilon}}g(z)=\sup_{z\in\partial U_{\operatorname{\Gamma},\varepsilon}}g(z),$
where $\partial U_{\operatorname{\Gamma},\varepsilon}$ denotes the boundary of the cuspidal neighborhood $U_{\operatorname{\Gamma},\varepsilon}$. Furthermore, from elementary calculus, we derive that
(20) $\sup_{z\in\mathbb{H}}h(z)=y^{2k}|q(z)|^{2}=y^{k}e^{-4\pi y}|_{y=k/2\pi}.$
Combining equation (2.2) with observations (17), (19), and (20), we infer that
$\sup_{z\in\operatorname{{\it{\overline{X}}}}}\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=\sup_{z\in\partial U_{\operatorname{\Gamma},k/2\pi}}\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=O_{\operatorname{\Gamma}}(k^{3/2}).$
which completes the proof of the proposition. ∎
In the following proposition, we ascertain the asymptotics of the Bergman kernel in the neighborhood of a cusp, for large values of $k$. Consequently, we derive a lower bound for the Bergman kernel.
Proposition 2.3.
(21) $\|\operatorname{\mathcal{B}_{\mathcal{L}}^{\it{k}}}(z)\|_{\mathrm{hyp}}\gg_{\operatorname{\Gamma}}k,$
where the implied constant depends only on the Fuchsian subgroup.
Proof.
For $k\gg 1$, and a fixed $\varepsilon>0$, and $z\in{\mathcal{F}}_{\operatorname{\Gamma},\varepsilon}$, with notation and arguments as in the proof of Proposition 2.2, we derive
$\|\operatorname{\mathcal{B}_{\mathcal{L}}^{\it{k}}}(z)\|_{\mathrm{hyp}}=\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=\frac{2k-1}{4\pi}-\Bigg{|}\frac{2k-1}{4\pi}\sum_{\gamma\in\operatorname{\Gamma}\backslash\{\mathrm{Id}\}}\frac{(4y)^{k}}{\big{(}z-\overline{\gamma z}\big{)}^{2k}}\cdot\frac{y^{k}}{\big{(}\overline{cz+d}\big{)}^{2k}}\bigg{|}\geq$
$\frac{2k-1}{4\pi}-\frac{2k-1}{4\pi}\bigg{(}\frac{16}{{\rm cosh}^{2k-4}(r_{\operatorname{\Gamma},\varepsilon}/4)}+\frac{8}{{\rm cosh}^{2k-3}(r_{\operatorname{\Gamma},\varepsilon}/2)}\bigg{)}-$
$\frac{2k-1}{2\pi{\rm sinh}^{2}(r_{\operatorname{\Gamma},\varepsilon}/4)}\bigg{(}\frac{1}{{\rm cosh}^{2k-3}(r_{\operatorname{\Gamma},\varepsilon}/2)}+\frac{1}{{\rm cosh}^{2k-4}(r_{\operatorname{\Gamma},\varepsilon}/2)}\bigg{)}\gg_{\operatorname{\Gamma}}k,$
(22) $\mathrm{where}\,\,r_{\operatorname{\Gamma},\varepsilon}:=\big{\{}\operatorname{{\it{d}}_{\mathrm{hyp}}}(z,\gamma z)\big{|}\,z\in{\mathcal{F}}_{\operatorname{\Gamma},\varepsilon},\,\mathrm{and}\,\gamma\in\operatorname{\Gamma}\backslash\{\mathrm{Id}\}\big{\}}.$
Furthermore, for a fixed $z=x+iy\in U_{\operatorname{\Gamma},\varepsilon}$, as $k$ approaches $\infty$, observe that absolute value of each of the term in the series
$\frac{2k-1}{4\pi}\sum_{\gamma\in\operatorname{\Gamma}_{i\infty}\backslash\{\mathrm{Id}\}}\frac{(2y)^{2k}}{\big{(}z-\overline{\gamma z}\big{)}^{2k}}=\frac{2k-1}{4\pi}\sum_{n\in{\mathbb{Z}}\backslash\{0\}}\frac{(2y)^{2k}}{\big{(}n+2iy\big{)}^{2k}}$
approach zero.
Hence, for $k\gg 1$, and $z\in U_{\operatorname{\Gamma},\varepsilon}$, we deduce that
$\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=$
$\frac{2k-1}{4\pi}-\frac{2k-1}{4\pi}\Bigg{|}\sum_{\gamma\in\operatorname{\Gamma}\backslash\operatorname{\Gamma}_{i\infty}}\frac{(4y)^{k}}{\big{(}z-\overline{\gamma z}\big{)}^{2k}}\cdot\frac{y^{k}}{\big{(}\overline{cz+d}\big{)}^{2k}}\Bigg{|}-\frac{2k-1}{4\pi}\Bigg{|}\sum_{n\in{\mathbb{Z}}\backslash\{0\}}\frac{(2y)^{2k}}{\big{(}n+2iy\big{)}^{2k}}\Bigg{|}=$
$\frac{2k-1}{4\pi}-\frac{2k-1}{4\pi}\bigg{(}\frac{16}{{\rm cosh}^{2k-4}(r_{\operatorname{\Gamma}}/4)}+\frac{8}{{\rm cosh}^{2k-3}(r_{\operatorname{\Gamma}}/2)}\bigg{)}-$
(23) $\frac{2k-1}{2\pi{\rm sinh}^{2}(r_{\operatorname{\Gamma}}/4)}\bigg{(}\frac{1}{{\rm cosh}^{2k-3}(r_{\operatorname{\Gamma}}/2)}+\frac{1}{{\rm cosh}^{2k-4}(r_{\operatorname{\Gamma}}/2)}\bigg{)}-\frac{2k-1}{4\pi}\Bigg{|}\sum_{n\in{\mathbb{Z}}\backslash\{0\}}\frac{(2y)^{2k}}{\big{(}n+2iy\big{)}^{2k}}\Bigg{|}\gg_{\operatorname{\Gamma}}k.$
Combining estimates (2.3) and (2.3), completes the proof of the proposition. ∎
@@ 2.3. Estimates of the Bergman metric
So employing the same arguments as in the proof of Proposition 2.3, we can conclude that
@@
Applying the above results, we now estimate the Bergman metric.
Theorem 2.7.
@@
$\sup_{z\in\operatorname{{\it{\overline{X}}}}}\bigg{|}\frac{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{Ber},{\it{k}}}}(z)}{\operatorname{\mu_{{\it\overline{X}}}^{\mathrm{hyp}}}(z)}\bigg{|}=O_{\operatorname{\Gamma}}(k^{3}),$
@@
For $k\geq 3$, and $z\in\operatorname{{\it{\overline{X}}}}\backslash\{i\infty\}$, from Lemma 2.4, we have
@@
We now estimate the second term on the right-hand side of the above inequality. For $z\in\operatorname{{\it{\overline{X}}}}$, from estimate (27), we derive
@@
Furthermore, for $k\gg 1$, from estimate (21), we have
(30) $\|\operatorname{\mathcal{B}_{{\Gamma}}^{2{\it{k}}}}(z)\|_{\mathrm{pet}}=\|\operatorname{\mathcal{B}_{\mathcal{L}}^{\it{k}}}(z)\|_{\mathrm{hyp}}\gg_{\operatorname{\Gamma}}k,$
@@
We now estimate the third term on the right-hand side of inequality (28). For $z\in\operatorname{{\it{\overline{X}}}}$, combining estimates (26) and (30), we arrive at the following estimate, for the third term on the right-hand side of inequality (28)
@@
Combining equation (28) with estimates (31) and (32), completes the proof of the theorem. ∎